# Further support for $S_{8}$ increasing with effective redshift

### v1

@@ Abstract
In Adil et al. (2023) Adil et al. (2023), we reported an increasing trend in $S_{8}$ with effective redshift $z_{\textrm{eff}}$ based on $f\sigma_{8}(z)$ constraints over the redshift range $0\lesssim z\lesssim 2$, and predicted that this trend would be observable in independent datasets. Recently, the studies by Artis et al. Artis et al. (2024) and the ACT+DESI collaboration Qu et al. (2024) appeared, presenting data that aligns with the expected trends. In this letter, we quantify the statistical significance of the increasing $S_{8}$ trends in Artis et al. (2024); Qu et al. (2024) by fitting a linear model to estimate the slope $\Delta\,S_{8}/\Delta\,z_{\textrm{eff}}$, and comparing the results to mock simulations. We find probabilities of $p=0.0163$ and $p=0.01893$, corresponding to approximately $2.1\sigma$ for each dataset. Using Fisher’s method to combine the independent probabilities, we obtain $p=0.0027$ ($2.8\sigma$). When we incorporate our earlier findings Adil et al. (2023), the combined statistical significance reaches between $3\sigma$ and $3.7\sigma$. This letter continues a series of studies initiated in 2020 that explore redshift-dependent $\Lambda$CDM parameters as an indication of a breakdown in the standard cosmological model.
@@ I Introduction
Naturally, such tests begin with $H_{0}$, a universal (integration) constant in any Friedmann-Lemaître-Robertson-Walker (FLRW) cosmology. The motivation is clear: the Hubble tension Aghanim et al. (2020); Riess et al. (2022); Freedman (2021); Pesce et al. (2020); Blakeslee et al. (2021); Kourkchi et al. (2020) suggests that $H_{0}$ may not be a constant under the assumption of the $\Lambda$CDM model. Corroborating support can be found in claims of descending trends in $H_{0}$ with effective redshift $z_{\rm eff}$ in the literature Wong et al. (2020); Millon et al. (2020); Krishnan et al. (2020); Dainotti et al. (2021, 2022); Ó Colgáin et al. (2022); Colgáin et al. (2024); Malekjani et al. (2024); Hu and Wang (2022); Jia et al. (2023); Vagnozzi (2023). Within $\Lambda$CDM, the matter density today ($z=0$), $\Omega_{m}$, is the next most relevant parameter in the late Universe (at low redshifts). Noting that the late Universe Hubble parameter $H(z)$ at $z\gtrsim 1$ is essentially governed by the combination $H_{0}^{2}\Omega_{m}$, one expects an increasing trend in $\Omega_{m}$ with effective redshift. This complementary signal has been reported in various forms Ó Colgáin et al. (2022); Colgáin et al. (2024); Malekjani et al. (2024); Ó Colgáin et al. (2024); Risaliti and Lusso (2019); Lusso et al. (2020); Yang et al. (2020); Khadka and Ratra (2020, 2021, 2022); Pourojaghi et al. (2022); Pastén and Cárdenas (2023)  [^1] Alternatively, if $\Omega_{m}$ is a genuine constant in line with the prevailing assumption, it is important to calibrate it model independently Pedrotti et al. (2024)..
@@
Besides the background parameters $H_{0},\Omega_{m}$, one may explore parameters describing perturbations, most notably the weighted amplitude of matter fluctuations $S_{8}:=\sigma_{8}\sqrt{\Omega_{m}/0.3}$, where a milder tension has persisted Aghanim et al. (2020); Heymans et al. (2013); Joudaki et al. (2017); Troxel et al. (2018); Hikage et al. (2019); Asgari et al. (2021); Abbott et al. (2022). Here too, if the model is breaking down, one does not expect $S_{8}$ to remain constant, especially since it explicitly depends on $\Omega_{m}$. In Adil et al. (2023), based on redshift-binned $f\sigma_{8}(z)$ constraints, it was demonstrated that $S_{8}$ increases with effective redshift. It is worth noting that White et al. (2022) reports an earlier increase of $\sigma_{8}$ with redshift for fixed $\Omega_{m}$ (see Fig. 11 of White et al. (2022)) [^2] Results are typically presented as $\sigma_{8}(z)$ versus $z$ White et al. (2022); Abbott et al. (2023); Bocquet et al. (2024). albeit with large uncertainties, in a cross-correlation of luminous red galaxies with CMB lensing, but attributes the effect to a statistical fluctuation. The purpose of this note is to reinforce the claim by demonstrating the trend in recent independent studies of SRG/eRosita number counts of massive halos Artis et al. (2024) and cross-correlations of CMB lensing with DESI Legacy imaging galaxies Qu et al. (2024), elucidating the combined statistical significance of these trends. Systematics aside, all observables point to a breakdown of the $\Lambda$CDM model in the linear regime. We note in passing that Esposito et al. (2022); Tutusaus et al. (2023); Manna and Desai (2024); Sailer et al. (2024) also indicate an increasing $S_{8}$ or $\sigma_{8}$ parameter with effective redshift, though we do not incorporate these trends. [^3] The anomaly in the late-time integrated Sachs-Wolfe effect in supervoids Kovács et al. (2019, 2022) also suggests weaker growth at lower and stronger growth at higher redshifts, thereby interpolating Planck expectations at intermediate redshifts. This weaker growth may also be interpreted in terms of the $\gamma$CDM model, e.g., Nguyen et al. (2023), but as remarked in Artis et al. (2024) with which we agree, the $\gamma$CDM model lacks an immediate physical interpretation. Retracing its origins to Wang and Steinhardt (1998), it is, at best, a simple and effective approximation.
@@ II Analysis
We now turn to the ACT DR6 Qu et al. (2023); Madhavacheril et al. (2023), Planck PR4 Carron et al. (2022), DESI Dey et al. (2019) constraints in Table 2, reproduced from Table 4 of Qu et al. (2024). The effective redshifts $z_{\textrm{eff}}\in\{0.37,\,0.51,\,0.65\}$ can be found in Table 1 of Qu et al. (2024). The average or “joint” constraint is $S_{8}=0.765$. The increasing trend of $S_{8}$ with $z_{\textrm{eff}}$ is evident in both Table 2 and Fig. 2. Note that, in contrast to Fig. 1, despite the studies disagreeing on whether the results are closer to Planck at lower or higher redshifts—suggesting potential systematics in the observables—the same increasing trend of $S_{8}$ with $z_{\textrm{eff}}$ remains evident. For completeness, we isolate the constraints from Table 4 of Qu et al. (2024) that correspond to Fig. 2. One could extend this analysis to include BAO data, but doing so would not further develop the point being made. Repeating the same steps as above, we find a slope of $m=0.59^{+0.29}_{-0.28}$ from MCMC analysis, which suggests a discrepancy of approximately $\sim 2.1\sigma$ from a horizontal line. Once again, we compare to mock simulations and find that the probability of obtaining a larger slope is $p=0.0185$ ($1851$ from $10^{5}$ realisations), corresponding to $\sim 2.1\sigma$ for a one-sided Gaussian, thus confirming the MCMC analysis. Given that the probabilities are independent, we can combine them using Fisher’s method to obtain $p=0.0027$. This confirms the increasing trend of $S_{8}$ with $z_{\textrm{eff}}$ across two independent data sets, with a statistical significance of approximately $2.8\sigma$.
@@
III Discussion

### v2

@@ Abstract
In Adil et al. (2023) Adil et al. (2023), we reported an increasing trend in $S_{8}$ with effective redshift $z_{\textrm{eff}}$ based on $f\sigma_{8}(z)$ constraints over the redshift range $0\lesssim z\lesssim 2$, and predicted that this trend would be observable in independent datasets. Recently, the studies by Artis et al. Artis et al. (2024) and the ACT/DESI collaboration Qu et al. (2024) appeared, presenting data that aligns with the expected trends. In this letter, we quantify the statistical significance of the increasing $S_{8}$ trends in Artis et al. (2024); Qu et al. (2024) by fitting a linear model to estimate the slope $\Delta\,S_{8}/\Delta\,z_{\textrm{eff}}$, and comparing the results to mock simulations. We find probabilities of $p=0.0163$ and $p=0.0185$, corresponding to approximately $2.1\sigma$ for each dataset. Using Fisher’s method to combine the independent probabilities, we obtain $p=0.0027$ ($2.8\sigma$). When we incorporate our earlier findings Adil et al. (2023), the combined statistical significance reaches between $3\sigma$ and $3.7\sigma$. Finally, we discuss how “scatter” in $\sigma_{8}$/$S_{8}$ constraints from recent DESI full-shape galaxy clustering fits this picture at low statistical significance. This letter continues a series of studies initiated in 2020 that explore redshift-dependent $\Lambda$CDM parameters as an indication of a breakdown in the standard cosmological model.
@@ I Introduction
Naturally, such tests begin with $H_{0}$, a universal (integration) constant in any Friedmann-Lemaître-Robertson-Walker (FLRW) cosmology. The motivation is clear: the Hubble tension Aghanim et al. (2020); Riess et al. (2022); Freedman (2021); Pesce et al. (2020); Blakeslee et al. (2021); Kourkchi et al. (2020) suggests that $H_{0}$ may not be a constant under the assumption of the $\Lambda$CDM model. Corroborating support can be found in claims of descending trends in $H_{0}$ with effective redshift $z_{\rm eff}$ in the literature Wong et al. (2020); Millon et al. (2020); Krishnan et al. (2020); Dainotti et al. (2021, 2022); Ó Colgáin et al. (2022); Colgáin et al. (2024); Malekjani et al. (2024); Hu and Wang (2022); Jia et al. (2023); Vagnozzi (2023). Within $\Lambda$CDM, the matter density today ($z=0$), $\Omega_{m}$, is the next most relevant parameter in the late Universe (at low redshifts). Noting that the late Universe Hubble parameter $H(z)$ at $z\gtrsim 1$ is essentially governed by the combination $H_{0}^{2}\Omega_{m}$, one expects an increasing trend in $\Omega_{m}$ with effective redshift. This complementary signal has been reported in various forms Ó Colgáin et al. (2022); Colgáin et al. (2024); Malekjani et al. (2024); Ó Colgáin et al. (2024); Risaliti and Lusso (2019); Lusso et al. (2020); Yang et al. (2020); Khadka and Ratra (2020, 2021, 2022); Pourojaghi et al. (2022); Pastén and Cárdenas (2023). [^1] Alternatively, if $\Omega_{m}$ is a genuine constant in line with the prevailing assumption, it is important to calibrate it model independently Pedrotti et al. (2024).
@@
Besides the background parameters $H_{0},\Omega_{m}$, one may explore parameters describing perturbations, most notably the weighted amplitude of matter fluctuations $S_{8}:=\sigma_{8}\sqrt{\Omega_{m}/0.3}$, where a milder tension has persisted Aghanim et al. (2020); Heymans et al. (2013); Joudaki et al. (2017); Troxel et al. (2018); Hikage et al. (2019); Asgari et al. (2021); Abbott et al. (2022). Here too, if the model is breaking down, one does not expect $S_{8}$ to remain constant, especially since it explicitly depends on $\Omega_{m}$. In Adil et al. (2023), based on redshift-binned $f\sigma_{8}(z)$ constraints, it was demonstrated that $S_{8}$ increases with effective redshift. It is worth noting that White et al. (2022) reports an earlier increase of $\sigma_{8}$ with redshift for fixed $\Omega_{m}$ (see Fig. 11 of White et al. (2022)) [^2] Results are typically presented as $\sigma_{8}(z)$ or $S_{8}(z)$ versus $z$ White et al. (2022); García-García et al. (2021); Abbott et al. (2023); Bocquet et al. (2024). albeit with large uncertainties, in a cross-correlation of luminous red galaxies with CMB lensing, but attributes the effect to a statistical fluctuation. The purpose of this note is to reinforce the claim by demonstrating the trend in recent independent studies of SRG/eROSITA number counts of massive halos Artis et al. (2024) and cross-correlations of CMB lensing with DESI Legacy imaging galaxies Qu et al. (2024), elucidating the combined statistical significance of these trends. Systematics aside, all observables point to a breakdown of the $\Lambda$CDM model in the linear regime. We note in passing that Esposito et al. (2022); Tutusaus et al. (2023); Manna and Desai (2024); Sailer et al. (2024) also indicate an increasing $S_{8}$ or $\sigma_{8}$ parameter with effective redshift, though we do not incorporate these trends. [^3] The anomaly in the late-time integrated Sachs-Wolfe effect in supervoids Kovács et al. (2019, 2022) also suggests weaker growth at lower and stronger growth at higher redshifts, thereby interpolating Planck expectations at intermediate redshifts. This weaker growth may also be interpreted in terms of the $\gamma$CDM model, e.g., Nguyen et al. (2023), but as remarked in Artis et al. (2024) with which we agree, the $\gamma$CDM model lacks an immediate physical interpretation. Retracing its origins to Wang and Steinhardt (1998), it is, at best, a simple and effective approximation.
@@ II Analysis
We now turn to the ACT DR6 Qu et al. (2023); Madhavacheril et al. (2023), Planck PR4 Carron et al. (2022), DESI Dey et al. (2019) constraints in Table 2, reproduced from Table 4 of Qu et al. (2024). The effective redshifts $z_{\textrm{eff}}\in\{0.37,\,0.51,\,0.65\}$ can be found in Table 1 of Qu et al. (2024). The average or “joint” constraint is $S_{8}=0.765$. The increasing trend of $S_{8}$ with $z_{\textrm{eff}}$ is evident in both Table 2 and Fig. 2. Note that, with reference to Fig. 1, despite the studies disagreeing on whether the results are closer to Planck at lower or higher redshifts—suggesting potential systematics in the observables—the same increasing trend of $S_{8}$ with $z_{\textrm{eff}}$ remains evident. For completeness, we isolate the constraints from Table 4 of Qu et al. (2024) that correspond to Fig. 2. One could extend this analysis to include BAO data, but doing so would not further develop the point being made. Repeating the same steps as above, we find a slope of $m=0.59^{+0.29}_{-0.28}$ from MCMC analysis, which suggests a discrepancy of approximately $\sim 2.1\sigma$ from a horizontal line. Once again, we compare to mock simulations and find that the probability of obtaining a larger slope is $p=0.0185$ ($1851$ from $10^{5}$ realisations), corresponding to $\sim 2.1\sigma$ for a one-sided Gaussian, thus confirming the MCMC analysis. Given that the probabilities are independent, we can combine them using Fisher’s method to obtain $p=0.0027$. This confirms the increasing trend of $S_{8}$ with $z_{\textrm{eff}}$ across two independent data sets, with a statistical significance of approximately $2.8\sigma$.
@@
III DESI Full-Shape Galaxy Clustering
DESI recently released cosmological constraints from the full-shape modelling of galaxy clusters Adame et al. (2024a, b). While the statistical significance of any signal for a $\Lambda$CDM deviation is low without external data, we explain how results are consistent with our claims. Given the overlap in DESI data with ACT/DESI observations Qu et al. (2024), one cannot combine $p$-values using Fisher’s method.
DESI analysed a bright galaxy sample (BGS) with redshift $0.40.1<z<$ 0.4, three luminous red galaxy (LRG) samples with redshifts $0.60.4<z<$ 0.6, $0.80.6<z<$ 0.8 and $1.10.8<z<$ 1.1, respectively, an emission line galaxy (ELG) sample in the range $1.61.1<z<$ 1.6 and a quasar (QSO) sample with $2.10.8<z<$ 2.1. In Table 3 we reproduce the effective redshift $z_{\textrm{eff}}$ from Table 1 Adame et al. (2024b) and the $\Omega_{m}$ and $\sigma_{8}$ constraints from Table 10 Adame et al. (2024b) for each tracer. We focus on full-modelling without BAO. As is evident from Table 3, the low redshift BGS $\Omega_{m}$ and $\sigma_{8}$ values are lower than the corresponding high redshift QSO values by $\sim 0.8\sigma$ and $\sim 2.2\sigma$, respectively. In contrast, the tracers probing the redshifts in between exhibit scatter about intermediate values with no definite trend.
Tracer Redshift ($z$) $\Omega_{m}$ $\sigma_{8}$ $S_{8}$
BGS $0.295$ $0.272\pm 0.027$ $0.664^{+0.098}_{-0.14}$ $0.63^{+0.12}_{-0.16}$
LRG1 $0.510$ $0.297\pm 0.021$ $0.844^{+0.083}_{-0.095}$ $0.84^{+0.11}_{-0.12}$
LRG2 $0.706$ $0.280\pm 0.020$ $0.888^{+0.071}_{-0.084}$ $0.858^{+0.096}_{-0.11}$
LRG3 $0.930$ $0.294\pm 0.025$ $0.810^{+0.069}_{-0.078}$ $0.802^{+0.099}_{-0.11}$
ELG2 $1.317$ $0.297^{+0.028}_{-0.036}$ $0.749^{+0.055}_{-0.064}$ $0.745^{+0.087}_{-0.10}$
QSO $1.491$ $0.310^{+0.030}_{-0.041}$ $0.950^{+0.070}_{-0.084}$ $0.97^{+0.11}_{-0.14}$
Table 3: Effective redshifts, $\Omega_{m}$, and $\sigma_{8}$ constraints reproduced from Table 1 and Table 10 of Adame et al. (2024b). We estimate the $S_{8}$ constraints as outlined in the text.
That being said, it is plausible that the $S_{8}=\sigma_{8}\sqrt{\Omega_{m}/0.3}$ values, which are omitted from Adame et al. (2024b), show a more definite trend, because $\Omega_{m}$ and $\sigma_{8}$ from the low redshift BGS sample are evidently both lower than the corresponding constraints from the QSO sample. See Ó Colgáin et al. (2024) for speculation of an increasing trend of $\Omega_{m}$ with $z_{\textrm{eff}}$ in DESI BAO constraints DESI Collaboration et al. (2024). Here we will estimate the $S_{8}$ constraints. To do so, we approximate the $\Omega_{m}$ and $\sigma_{8}$ constraints as antisymmetric Gaussian distributions. More precisely, we assume that the distributions for values larger and smaller than the median ( $50^{\textrm{th}}$ percentile) are Gaussian with standard deviations dictated by the upper and lower errors, respectively. In the case of symmetric errors, the distribution reduces to a standard Gaussian. For each “half” of the Gaussian, we randomly generate $2\times 10^{5}$ odd configurations in a normal distribution, which we restrict to exactly $1.5\times 10^{5}$ configurations before combining to get $3\times 10^{5}$ configurations in the full distribution. By construction, this leads to $16^{\textrm{th}},50^{\textrm{th}}$ and $84^{\textrm{th}}$ percentiles that recover the central values and the errors up to small numbers for both $\Omega_{m}$ and $\sigma_{8}$. Given that our distribution for $\Omega_{m}$ and $\sigma_{8}$ has the same number of configurations, we combine them to get an $S_{8}$ distribution and extract the $16^{\textrm{th}},50^{\textrm{th}}$ and $84^{\textrm{th}}$ to record them in Table 3, thus providing us an estimate of the $S_{8}$ constraints from each tracer. The $S_{8}$ values appear against $z_{\textrm{eff}}$ in Fig. 3.
Figure 3: $S_{8}$ versus effective redshift for DESI full-shape modelling of galaxy clusters.
As is evident from Fig. 3, if there is an increasing trend of $S_{8}$ with $z_{\textrm{eff}}$, it is not significant. At this juncture, we assume the errors are Gaussian by adopting the largest error and repeat the earlier analysis. From MCMC analysis, we find a slope of $m=0.07^{+0.13}_{-0.12}$ for all tracers and a slope of $m=0.50^{+0.46}_{-0.45}$ for the first three data points in Fig. 3, where one should note that these constraints overlap in redshift range with the SRG/eROSITA and ACT/DESI observations. This motivates restricting the redshift range. Regardless, the statistical significance is low at $\sim 0.6\sigma$ to $\sim 1.1\sigma$ in DESI full-shape results Adame et al. (2024a, b). We further remark that the BGS and QSO $S_{8}$ values now differ at $\sim 1.8\sigma$. In switching from $\sigma_{8}$ to $S_{8}$, the errors inflate, but this effect is countered by some movement in the $\Omega_{m}$ central values. [^5] Under similar assumptions, replacing $S_{8}$ with $\sigma_{8}$ marginally decreases the statistical significance of an increasing trend with $z_{\textrm{eff}}$ $\sim 0.6\sigma\rightarrow\sim 0.5\sigma$ for all redshifts and marginally increases the statistical significance of the trend $\sim 1.1\sigma\rightarrow\sim 1.3\sigma$ if one restricts to the lowest three effective redshifts.
IV Discussion
@@
Our work departs from reconstructions of $S_{8}(z)/\sigma_{8}(z)$ White et al. (2022); García-García et al. (2021); Abbott et al. (2023); Bocquet et al. (2024) in two key ways. First, while $S_{8}(z)/\sigma_{8}(z)$ reconstructions can show deviations from Planck-$\Lambda$CDM behaviour, and identify redshift ranges where physics is missing, one still needs to separate missing scale-dependent physics from missing redshift-dependent physics. Here, this is less of a concern as ACT/DESI cross-correlations, SRG/eROSITA observations and $f\sigma_{8}(z)$ constraints from redshift space distortions typically only probe linear scales. Secondly, $S_{8}(z)/\sigma_{8}(z)$ reconstructions fail to map the problem back into constant $\Lambda$CDM fitting parameters showing that the $\Lambda$CDM model has broken down. With $S_{8}(z)/\sigma_{8}(z)$ reconstructions one has the weaker statement that the $\Lambda$CDM model is not tracking the data, not the stronger statement that the $\Lambda$CDM model fails to make unique predictions.